OpenStudy (anonymous):

12 squared (-6) to the 12th power

OpenStudy (saifoo.khan):

can you please write that more clearly, using the Equation editor?

OpenStudy (anonymous):

\[12\sqrt{(-6)^{12}}\]

OpenStudy (anonymous):

sorrry didnt know this was there

OpenStudy (saifoo.khan):

\[\Huge \sqrt[12]{-6^{12}}\] OR \[\Huge 12\sqrt{-6^{12}}\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

-6

OpenStudy (saifoo.khan):

6

OpenStudy (saifoo.khan):

Answer.

OpenStudy (anonymous):

can you explain to me how to get this answer please

OpenStudy (anonymous):

it is \[(-6)^{\left(\begin{matrix}12 \\ 12\end{matrix}\right)}\]

OpenStudy (saifoo.khan):

\[\Huge -6^\frac{12}{12}\]

OpenStudy (anonymous):

the 12 at the bottom is the root while the 12 at the uppper is the power so when you devide both the onswer will be raise to one

OpenStudy (anonymous):

oh ok ty

OpenStudy (saifoo.khan):

please hit good answer for both.

OpenStudy (anonymous):

I have never been on here before it is nice ty

OpenStudy (saifoo.khan):

No problem!